# AdvRepair: Provable Repair of Adversarial Attack

### v1

@@ 3.2 Training the patch modules
The first challenge is to train the patch modules, which are each implemented as a fully connected neural network, to repair each robustness property $(N,B(x_{i},r))$ with a provable guarantee. To achieve provability, we employ formal verification in training these patch modules. The result of verification on a robustness property can be transformed into a loss function; once this loss reaches zero, the property is verified to be true, and the patch module in this stage can defend against all possible adversarial attacks in this robustness region. This idea has been widely adopted in DNN repair [21, 39].
@@
In this work, we employ DeepPoly [58] as the verification engine. DeepPoly uses abstract interpretation to give every neuron an upper/lower bound in the form of an affine function, where only variables in previous layers occur, and the numerical bound can be derived by propagating backwards these affine upper/lower bounds to the input layer. Recall that a local robustness property $(N,B(x_{i},r))$ holds iff for any $x\in B(x_{i},r)$ and $\ell\neq\ell_{0}=C_{N}(x_{i})$, $N(x)_{\ell}<N(x)_{\ell_{0}}$, so DeepPoly calculates the abstraction of $N$ on $B(x_{i},r)$ for every neuron and the expressions $N(x)_{\ell}-N(x)_{\ell_{0}}$ for $\ell\neq\ell_{0}$. From the abstraction, we can obtain an affine function in which there are only input variables as a sound upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$, i.e.,
$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},$ (1)
@@
where $\alpha_{\ell}\in\mathbb{R}^{n_{0}}$ and $\beta_{\ell}\in\mathbb{R}$ are constants. It is easy to obtain the numerical upper bound of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on the box region $B(x_{i},r)$. If this upper bound is negative for every $\ell\neq\ell_{0}$, then the local robustness property $(N,B(x_{i},r))$ is verified to be true by DeepPoly. Therefore, it is natural to use this upper bound as the loss function to train the corresponding patch module.
Definition 3
@@
For a local robustness property $\varphi=(N,B(x_{i},r))$, we define the safety violated loss function of $\varphi$ as
@@
where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is the upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$ given by DeepPoly, as shown in Eq. (1). For local robustness properties $\varphi_{1},\ldots,\varphi_{k}$ which share the same ground truth label $\ell_{0}$, we define $\mathcal{L}(\bigwedge_{i=1}^{k}\varphi_{i})=\sum_{i=1}^{k}\mathcal{L}(\varphi_{i})$.
@@
For a local robustness property $\varphi=(N,B(x_{i},r))$, its safety violated loss function $\mathcal{L}(\varphi)$ being 00 implying that $\varphi$ is verified to be true by DeepPoly.
Theorem 3.1
Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
@@
where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
When we train a patch module $P$ on the DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$, the maximum of the safety violated loss function of a local robustness property $\varphi$, namely $\mathcal{L}^{*}(\varphi)$, is parameterized with the weights $w$ in $P$, and this $\mathcal{L}^{*}(\varphi)(w)$ is the loss function for training this patch $P$. By employing the standard gradient descent method, we can train the patch $P$, until the loss reaches 00 or the number of epochs reaches a threshold $R$. We claim that, even if $\mathcal{L}^{*}(\varphi)(w)$ does not reach 00, the repair does not necessarily fail, since the upper bound $\mathcal{L}^{*}(\varphi)$ given by DeepPoly is not exact; as it decreases in the training process, $F$ is growing more likely to satisfy $\varphi$. Also, to improve the precision of $\mathcal{L}^{*}(\varphi)$ obtained from DeepPoly, we employ the input interval partitioning technique from ART [39]. It selects the partition dimension by computing the multiplication of the partial derivative of the safety violated loss function $\mathcal{L}(\varphi)(x)$ and the size of the input interval in the corresponding dimension, and bisects the box region over the dimension with the maximum score.
Algorithm 1 AdvRepair (Property-guided)
1:
2:Original DNN $N$, pairs of input and its adversarial examples $\{(x_{i},x_{i}^{*})\}_{i=1}^{n}$, and radius $r>0$
3:
4:A repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$
5:$\mathcal{C}\leftarrow(B(x_{1},r),\ldots,B(x_{n},r))$, $\mathrm{iter}\leftarrow 0$
@@
6:$\mathrm{Init}(\mathcal{P})$triangleright Initialize the patch modules $\mathcal{P}=(P_{1},\ldots,P_{n})$
@@
7:$D[\cdot,\cdot]\leftarrow\{\}$triangleright A dictionary where $D[i]$ stores properties fixed with $P_{i}$
8:for $i\leftarrow 1~{}\text{to}~{}n$ do
9:     $D[i]\leftarrow\{(F,B(x_{i},r))\}$
10:$E\leftarrow\{1,\ldots,m\}$triangleright Record the properties not fixed yet
11:while $\mathrm{iter}<M$ do $\triangleright$ $M$: Maximum number of iterations
12:     $\mathrm{iter}\leftarrow\mathrm{iter}+1$
13:     for $j\in E$ do
14:         $(P_{j},T,\mathrm{Repaired})\leftarrow\textsc{Train}(N,P_{j},D[j])$triangleright Alg. 2
15:         if $\mathrm{Repaired}=\mathbf{True}$ then
16:              $E\leftarrow E\setminus\{j\}$
17:         else
18:              for $\varphi\in T$ do $\triangleright$ We write $\varphi=(F,X)$
19:                  for $d\leftarrow 1~{}\text{to}~{}n_{0}$ do
@@
20:                       $\operatorname{score}_{d}\leftarrow\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|$
21:                  $d^{*}\leftarrow\arg\max_{d}\operatorname{score}_{d}$
22:                  $X_{1},X_{2}\leftarrow\mathrm{Bisect}(X,d^{*})$triangleright Bisection on the $d^{*}$th dimension
23:                  $\varphi_{1}\leftarrow(F,X_{1})$, $\varphi_{2}\leftarrow(F,X_{2})$, $D[j]\leftarrow(D[j]\cup\{\varphi_{1},\varphi_{2}\})\setminus\{\varphi\}$
24:     if $E=\{\}$ then
25:         return $F$triangleright Repair with provable guarantee
26:return $F$triangleright Repair without provable guarantee
@@
1:
2:Original DNN $N$, DNN $P$ as a patch module, and a finite set $U$ of local robustness properties
3:
4:The optimized patch module $P$, the set $T\subseteq U$ of properties to be refined, and whether the properties in $U$ have all been fixed with a provable guarantee
5:function Train($N,P,U$)
6:     $\mathrm{epoch}\leftarrow 0$
7:     while $\mathrm{epoch}<R$  do $\triangleright$ $R$: maximum number of epochs
8:         $\mathrm{epoch}\leftarrow\mathrm{epoch}+1$
9:         $w\leftarrow w-\eta\cdot\nabla\mathcal{L}^{*}(\bigwedge U)(w)$triangleright $w$: the weights in $P$
10:         if $\mathcal{L}^{*}(\bigwedge U)(w)=0$ then $\triangleright$ $\mathcal{L}^{*}(\bigwedge U):=\mathcal{L}^{*}(\bigwedge_{\varphi\in U}\varphi)=\sum_{\varphi\in U}\mathcal{L}^{*}(\varphi)$
11:              return $(P,\{\},\mathbf{True})$
12:     $T\leftarrow\mathrm{Slice}_{K}(\mathrm{ArgSort}_{\varphi}\{\mathcal{L}^{*}(\varphi)(w)\mid\mathcal{L}^{*}(\varphi)(w)>0,\varphi\in U\})$
13:     return $(P,T,\mathbf{False})$
The main algorithm of AdvRepair is shown in Alg. 1. We construct a dictionary $D$, where $D[i]$, initialized as $\{(F,B(x_{i},r))\}$, stores the properties that the $i$th patch module $P_{i}$ repairs (Line 3–5, Alg. 1). The repair process has at most $M$ iterations. In each iteration, we train the patches which does not yet provide provable repair, i.e., those whose index is in $E$. The algorithm of training a specific patch module $P$ for repairing a set $U$ of properties is shown in Alg. 2, which outputs the optimized patch module $P$, the set $T$ of properties to be refined, and a boolean label recording whether the repair of the patch $P$ has a provable guarantee. A standard gradient descent procedure is run until a provable repair is achieved, i.e., $\mathcal{L}^{*}(\bigwedge U)(w)=0$, or the number of epochs reaches a threshold $R$ (Line 3–7, Alg. 2). After that, if it is still not provable, we sort $\mathcal{L}^{*}(\varphi)(w)$ for $\varphi\in U$ from the largest to the smallest, and extract the largest $K$ ones (whose value is strictly larger than 00) as the properties to be refined (Line 8, Alg. 2). For a property $\varphi=(F,X)$ to be refined, we select an input dimension to bisect the input space $X$. For an input dimension $d$, we define
@@
where $\partial_{d}\mathcal{L}(\varphi)(x)$ is the partial derivative of $\mathcal{L}(\varphi)(x)$ on the $d$th dimension. We choose the dimension with the largest score to bisect $X$, and the property $\varphi=(F,X)$ is refined to two properties $\varphi_{1}=(F,X_{1})$ and $\varphi_{2}=(F,X_{2})$, recorded in the dictionary $D$ (Line 14–19, Alg. 1). At the end of each iteration, we check whether all the patches provide provable repair; if so, it terminates immediately and outputs the current repaired DNN $F$, and this repair is provable (Line 20–21, Alg. 1). If $E$ is still non-empty after $M$ iterations, it outputs $F$ without provable guarantee (Line 22, Alg. 1). Although we are focusing on fixing adversarial attacks in this work, AdvRepair also works for repair safety properties of DNNs by making minor changes in Alg. 1.
@@
In AdvRepair, we follow a classical way of defining the loss function with the linear relaxation obtained from DeepPoly, and incorporate it into our patch-based repair framework. This combination has several advantages. First, every patch is responsible for repairing properties in a specific pattern, which makes it much easier to obtain a provable repair; even if the repair is not provable, the safety violated loss is significantly declining in the training process, and it is highly possible that the repaired DNN is locally robust. Different from the neuron-level repair, there is no modification on the original DNN $N$ in AdvRepair, and due to the design of the indicator, the patch modules does not affect the behaviors of other input before $\tau$ allocates a patch to it. This mechanism avoids the drawdown of accuracy, which is quite severe in many neuron-level repair methods. Also, by allocating patches with $\tau$, we can achieve good generalization of our repair to other inputs, and it is the key to solving this essential challenge in fixing adversarial attacks. In the following, we will present how to construct the patch allocation function $\tau$ to improve generalization.
3.3 Patch allocation
For the property-guided method in AdvRepair, we need to define a patch allocation function $\tau$ because the construction of patches in this way does not naturally generalize to other inputs. This patch allocation function $\tau$ is the key to achieving good generalization of the patch modules trained locally to global inputs.
To match the best patch modules for an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we need to extract some specific patterns of the patch modules which generalize to other inputs and their adversarial robustness. Such patterns may include:
  • the ground truth of $x^{\prime}$, which can be represented with $\ell_{0}=C_{N}(x^{\prime})$,
  • the targeted label, i.e., the index of the second largest entry in $N(x^{\prime})$,
[…]
In the implementation of $\tau$, the patches with the same ground truth label as $x^{\prime}$ is alternative for $x^{\prime}$, and we allocate the ten with the maximum cosine similarity between $N(x^{\prime})$ and $N(x_{i})$ to $x^{\prime}$. Here we do not match with the targeted label because it is possible that more than one labels can potentially attack $N$ in the robustness region $B(x^{\prime},r)$, and matching by the cosine similarity can cover this possibility.

### v2

@@ III-B Training the patch modules
The first challenge is to train the patch modules, which are each implemented as a fully connected neural network, to repair each robustness property $(F,B(x_{i},r))$ with a provable guarantee. To achieve provability, we employ formal verification in training these patch modules. The result of verification on a robustness property can be transformed into a loss function; once this loss reaches zero, the property is verified to be true, and the patch module in this stage can defend against all possible adversarial attacks in this robustness region. This idea has been widely adopted in DNN repair [24, 25].
@@
In this work, we employ DeepPoly [23] as the verification engine [^1] Also there are several other verification tools based on abstract interpretation, such as CROWN [26] and DeepZ [27], which vary in precision, affecting loss function evaluations. We choose DeepPoly since it provides a good balance between efficiency and precision.. DeepPoly uses abstract interpretation to give every neuron an upper/lower bound in the form of an affine function, where only variables in previous layers occur, and the numerical bound can be derived by propagating backwards these affine upper/lower bounds to the input layer. Recall that a local robustness property $(F,B(x_{i},r))$ holds iff for any $x\in B(x_{i},r)$ and $\ell\neq\ell_{0}=C_{F}(x_{i})$, $F(x)_{\ell}<F(x)_{\ell_{0}}$, so DeepPoly calculates the abstraction of $F$ on $B(x_{i},r)$ for every neuron and the expressions $F(x)_{\ell}-F(x)_{\ell_{0}}$ for $\ell\neq\ell_{0}$. From the abstraction, we can obtain an affine function in which there are only input variables as a sound upper bound of $F(x)_{\ell}-F(x)_{\ell_{0}}$ on $B(x_{i},r)$, i.e.,
$\forall x\in B(x_{i},r),F(x)_{\ell}-F(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},$ (1)
@@
where $\alpha_{\ell}\in\mathbb{R}^{n_{0}}$ and $\beta_{\ell}\in\mathbb{R}$ are constants. It is easy to obtain the numerical upper bound of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on the box region $B(x_{i},r)$. If this upper bound is negative for every $\ell\neq\ell_{0}$, then the local robustness property $(F,B(x_{i},r))$ is verified to be true by DeepPoly. Therefore, it is natural to use this upper bound as the loss function to train the corresponding patch module.
Definition 4.
@@
For a local robustness property $\varphi=(F,B(x_{i},r))$, we define the safety violated loss function as
@@
where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is the upper bound of $F(x)_{\ell}-F(x)_{\ell_{0}}$ on $B(x_{i},r)$ given by DeepPoly, as shown in Eq. (1). For local robustness properties $\varphi_{1},\ldots,\varphi_{k}$ which share the same ground truth label $\ell_{0}$, we define $\mathcal{L}(\bigwedge_{i=1}^{k}\varphi_{i})=\sum_{i=1}^{k}\mathcal{L}(\varphi_{i})$.
@@
For a local robustness property $\varphi=(F,B(x_{i},r))$, its safety violated loss function $\mathcal{L}(\varphi)$ being 00 implies that $\varphi$ is verified to be true by DeepPoly.
Theorem 1.
Let $\varphi=(F,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=$ $\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot$
$(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
@@
where $\mathrm{elmax}$ and $\mathrm{elmin}$ are the element-wise $\max$ and $\min$ operation, $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
The notion that $\mathcal{L}^{*}(\varphi)$ is the expansion of $\mathcal{L}(\varphi)(x)$ after taking corresponding values on the boundary of the ball. To improve the precision of $\mathcal{L}^{*}(\varphi)$ obtained from DeepPoly, we employ the input interval partitioning technique from ART [25]. It selects the partition dimension by computing the multiplication of the partial derivative of the safety violated loss function $\mathcal{L}(\varphi)(x)$ and the size of the input interval in the corresponding dimension, and bisects the box region over the dimension with the maximum score. After partitioning, the property $\varphi$ is split into two new properties, whose input sets are two sub-boxes of the original input region, and the union of these two sub-boxes is the input set of $\varphi$.
Algorithm 1 PatchPro
1:Original DNN $N$, pairs of input and its adversarial examples $\{(x_{i},x_{i}^{*})\}_{i=1}^{n}$, and radius $r>0$
2:A repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$
3:$\mathcal{C}\leftarrow(B(x_{1},r),\ldots,B(x_{n},r))$, $\mathrm{iter}\leftarrow 0$
@@
4:$\mathrm{Init}(\mathcal{P})$triangleright Initialize the patch modules $\mathcal{P}=(P_{1},\ldots,P_{n})$
@@
5:$D[\cdot,\cdot]\leftarrow\{\}$triangleright A dictionary where $D[i]$ stores properties fixed with $P_{i}$
6:for $i\leftarrow 1~{}\text{to}~{}n$ do
7:    $D[i]\leftarrow\{(F,B(x_{i},r))\}$
8:$E\leftarrow\{1,\ldots,n\}$triangleright Record the properties not fixed yet
9:while $\mathrm{iter}<M$ do $\triangleright$ $M$: Maximum number of iterations
10:    $\mathrm{iter}\leftarrow\mathrm{iter}+1$
11:    for $j\in E$ do
12:        $(P_{j},T,\mathrm{Repaired})\leftarrow\textsc{Train}(N,P_{j},D[j])$triangleright Alg. 2
13:        if $\mathrm{Repaired}=\mathbf{True}$ then
14:           $E\leftarrow E\setminus\{j\}$
15:        else
16:           for $\varphi\in T$ do $\triangleright$ We write $\varphi=(F,X)$
17:               for $d\leftarrow 1~{}\text{to}~{}n_{0}$ do
@@
18:                  $\operatorname{score}_{d}\leftarrow\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|$
19:               $d^{*}\leftarrow\arg\max_{d}\operatorname{score}_{d}$
20:               $X_{1},X_{2}\leftarrow\mathrm{Bisect}(X,d^{*})$triangleright Bisection on the $d^{*}$th dimension
21:               $\varphi_{1}\leftarrow(F,X_{1})$, $\varphi_{2}\leftarrow(F,X_{2})$, $D[j]\leftarrow(D[j]\cup\{\varphi_{1},\varphi_{2}\})\setminus\{\varphi\}$
22:    if $E=\{\}$ then
23:        return $F$triangleright Repair with provable guarantee
24:return $F$triangleright Repair without provable guarantee
@@
1:Original DNN $N$, DNN $P$ as a patch module, and a finite set $U$ of local robustness properties
2:The optimized patch module $P$, the set $T\subseteq U$ of properties to be refined, and whether the properties in $U$ have all been fixed with a provable guarantee
3:function Train($N,P,U$)
4:    $\mathrm{epoch}\leftarrow 0$
5:    while $\mathrm{epoch}<R$  do $\triangleright$ $R$: maximum number of epochs
6:        $\mathrm{epoch}\leftarrow\mathrm{epoch}+1$
7:        $w\leftarrow w-\eta\cdot\nabla\mathcal{L}^{*}(\bigwedge U)(w)$triangleright $w$: the weights in $P$
8:        if $\mathcal{L}^{*}(\bigwedge U)(w)=0$ then $\triangleright$ $\mathcal{L}^{*}(\bigwedge U):=\sum_{\varphi\in U}\mathcal{L}^{*}(\varphi)$
9:           return $(P,\{\},\mathbf{True})$
10:    $T\leftarrow\mathrm{Slice}_{K}(\mathrm{ArgSort}_{\varphi}\{\mathcal{L}^{*}(\varphi)(w)\mid\mathcal{L}^{*}(\varphi)(w)>0,\varphi\in U\})$
11:    return $(P,T,\mathbf{False})$
The main algorithm of PatchPro is shown in Alg. 1. We construct a dictionary $D$, where $D[i]$, initialized as $\{(F,B(x_{i},r))\}$, stores the properties that the $i$th patch module $P_{i}$ repairs (Line 3–5, Alg. 1). The repair process has at most $M$ iterations. In each iteration, we train the patches which do not yet provide provable repair, i.e., those whose index is in $E$. The algorithm of training a specific patch module $P$ for repairing a set $U$ of properties is shown in Alg. 2, which outputs the optimized patch module $P$, the set $T$ of properties to be refined, and a boolean label recording whether the repair of the patch $P$ has a provable guarantee. A standard gradient descent procedure is run until a provable repair is achieved, i.e., $\mathcal{L}^{*}(\bigwedge U)(w)=0$, or the number of epochs reaches a threshold $R$ (Line 3–7, Alg. 2). After that, if it is still not provable, we sort $\mathcal{L}^{*}(\varphi)(w)$ for $\varphi\in U$ from the largest to the smallest, and extract the largest $K$ ones (whose value is strictly larger than 00) as the properties to be refined (Line 8, Alg. 2). For a property $\varphi=(F,X)$ to be refined, we select an input dimension to bisect the input space $X$. For an input dimension $d$, we define
@@
where $\partial_{d}\mathcal{L}(\varphi)(x)$ is the partial derivative of $\mathcal{L}(\varphi)(x)$ on the $d$th dimension. We choose the dimension with the largest score to bisect $X$, and the property $\varphi=(F,X)$ is refined to two properties $\varphi_{1}=(F,X_{1})$ and $\varphi_{2}=(F,X_{2})$, recorded in the dictionary $D$ (Line 14–19, Alg. 1). At the end of each iteration, we check whether all the patches provide provable repair; if so, it terminates immediately and outputs the current repaired DNN $F$, and this repair is provable (Line 20–21, Alg. 1). If $E$ is still non-empty after $M$ iterations, it outputs $F$ without provable guarantee (Line 22, Alg. 1). Although we are focusing on fixing adversarial attacks in this work, PatchPro also works for repair safety properties of DNNs by making minor changes in Alg. 1.
@@
In PatchPro, we follow a classical way of defining the loss function with the linear relaxation obtained from DeepPoly, and incorporate it into our patch-based repair framework. This combination has several advantages. First, every patch is responsible for repairing properties in a specific pattern, which makes it much easier to obtain a provable repair; even if the repair is not provable, the safety violated loss is significantly declining in the training process, and it is highly possible that the repaired DNN is locally robust. Different from the neuron-level repair, there is no modification on the original DNN $N$ in PatchPro, and due to the design of the indicator, the patch modules do not affect the behaviors of other input before $\tau$ allocates a patch to it. This mechanism avoids the drawdown of accuracy, which is quite severe in many neuron-level repair methods. Also, by allocating patches with $\tau$, we can achieve good generalization of our repair to other inputs, and it is the key to solving this essential challenge in fixing adversarial attacks. In the following, we will present how to construct the patch allocation function $\tau$ to improve generalization.
For a network $N$ with $L$ layers and a maximum of $n$ neurons per layer, the time complexity of running Deeppoly for one neuron is $O(n^{2}\cdot L)$ [23]. The time complexity of running Algorithm 2 once, which involves backpropagation for gradient computation and update weights of the patch networks in line 5 of this algorithm, is $O(R\cdot n_{\mathrm{patch}}\cdot L_{\mathrm{patch}})$, where $n_{\mathrm{patch}}$ is the maximum number of neurons per layer in all patch networks, and $L_{\mathrm{patch}}$ is the number of layers in the patch networks. Furthermore, during the execution of PatchPro, difficult-to-repair properties are split into two easier-to-repair properties, which may increase the number of properties to repair up to a predefined upper limit, denoted as $K$. In summary, the overall time complexity of PatchPro is $O(R\cdot M\cdot K\cdot(n^{3}\cdot L^{2}+n_{\mathrm{patch}}^{3}\cdot L_{\mathrm{patch}}^{2}))$.
III-C Patch allocation
For generalization purposes, it is necessary to introduce a patch allocation function $\tau$, given that the current patch construction approach lacks inherent adaptability to in-distribution data points. This patch allocation function $\tau$ is the key to achieving good generalization of the patch modules trained locally to global inputs.
To match the best patch modules for an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we propose utilizing the prediction $\ell_{0}=C_{N}(x^{\prime})$ of the original network on $x^{\prime}$ as a guiding principle, with the aim of selecting patches that share the same ground truth label as $\ell_{0}$. Accordingly, we formally define the set of patch modules $\tau(x^{\prime})$ associated with input $x^{\prime}$ as $\tau(x^{\prime})=\{i\mid C_{N}(x^{\prime})=C_{N}(x_{i})\}$. This definition establishes that the set $\tau(x^{\prime})$ encompasses all indices $i$ satisfying the condition $C_{N}(x^{\prime})=C_{N}(x_{i})$, i.e., those instances where the prediction $\ell_{0}$ made by the original network on $x^{\prime}$ is identical to the prediction made on sample $x_{i}$ at index $i$. By determining the patch module set $\tau(x^{\prime})$ in this manner, we ensure consistency between the selected modules and the input $x^{\prime}$ in terms of their predictions by the original network. This, in turn, is expected to enhance the robustness repair effectiveness for the specific input $x^{\prime}$.
Once a set of patch modules has been allocated to an input $x^{\prime}$, these modules effectively repair the entire robustness region of that input. Namely, we establish a new local robustness property such that any subsequent inputs $x\in B(x^{\prime},r)$ will also utilize the same set of patch modules as $x^{\prime}$, with the system’s response given by $F(x)=\sum_{j\in\tau(x^{\prime})}P_{j}(x)+N(x)$. In this context, the defense mechanism is established prior to the occurrence of adversarial attacks, obviating the need to re-allocate patch modules for each individual input $x$ within the robustness region, which would otherwise require computing $\tau(x)$ separately. This proactive construction of the defense ensures a consistent and efficient protection strategy against potential adversarial threats across the entire neighborhood of $x^{\prime}$, reinforcing the overall robustness.